OpenStudy (anonymous):

Finding the limit algebraically? Medal award. lim x^2 - 16 / x + 4 x -> -4 Answer choices : 1 -8 does not exist -4 I plugged in -4 and got zero for the numerator and denominator...so I chose "does not exist" But that answer is incorrect. What am I doing wrong? Thanks for any help!!

sam (.sam.):

There is a solution :)

sam (.sam.):

Try factoring the numerator

OpenStudy (anonymous):

okay! so (x -4)(x + 4)

sam (.sam.):

\[\lim_{x \rightarrow -4} \frac{(x+4)(x-4)}{x+4} \\ \\ \lim_{x \rightarrow -4} \frac{\cancel{(x+4)}(x-4)}{\cancel{(x+4)}} \\ \\ \lim_{x \rightarrow -4}x-4\] now substitute x=-4 into that

OpenStudy (anonymous):

oh!! so the answer would be -8!

sam (.sam.):

Yes!

OpenStudy (anonymous):

Thanks so much!! Never thought of that!

OpenStudy (anonymous):

@.Sam. nice use of \(\cancel{a}\)

sam (.sam.):

most of the limit problems involves factoring and playing with the terms, you just have to look at it carefully first :)